a flexible chain oh weight W hangs between 2 fixed points A & B which are at same horizontal level the inclination of the chain with the horizontal is x what is the tension of the chain at the mid point
 

Grade:11

2 Answers

Vikas TU
4 years ago
Hii 
Appl the tension at one end and midpoint , to draw the free body diagram of The Tension in chain.
T2 = T1 cosQ 
Where T2 is tension at midpoint ,and T1 is te tensin at end point.
T1 = M/2 * gsinQ 
T2 = Mg/2 * cot Q
Khimraj
4 years ago
Balance the forces to get the answer. 
T2 = T1* Cosθ 
T1 = M/2 * g * Sinθ 

So T2 = Mg/2 * Cotθ.
........................................................................